mathslover (mathslover):

Prove that : \(\large{^n P_r = ^{n-1} P_r + r ^{n-1} P _{r-1} }\)

mathslover (mathslover):

\(\large{^n P_r = ^{n-1} P_r + r ^{n-1} P _{r-1} }\)

mathslover (mathslover):

Sorry for that mistake. Please refresh the page to see the edited question or just check the above post by me.

mathslover (mathslover):

Here is my work : \( \large{ \textbf{Simplifying LHS to get RHS } \\ ^{n-1} P _ r + r ^{n-1} P _ {r-1} = \cfrac{(n-1)!}{(n-1-r)!} + r [ \cfrac{(n-1)!}{(n-r)!} ] \\ \cfrac{(n-1)! (n-r)}{(n-r-1)! (n-r)} + r [ \cfrac{(n-1)!}{(n-r)!} ] \\ \cfrac{(n-1)! (n-r)}{(n-r)!} + r [ \cfrac{(n-1)!}{(n-r)!} ] \\ \cfrac{(n-1)!}{(n-r)!} \{ n-r + r\} \\ \cfrac{(n-1)! n}{(n-r)!} \\ \cfrac{n!}{(n-r)!} = ^n P _r ... \mathsf{Proved!} \\ \\ ^{n-1} P _ r + r ^{n-1} P _ {r-1} = \cfrac{n!}{(n-r)!} = ^n P _ r } \) WOW I did it :)

mathslover (mathslover):

:O Well, OS helps in this way too ... Never thought about that.

OpenStudy (ajprincess):

Well done:)

mathslover (mathslover):

Thanks @ajprincess ... Actually I was not able to find any way to proceed for the proof. But I did it finally :) btw, you were typing something, was it the solution? Was it the same as mine or something different? If diff. then please let me know.

OpenStudy (ajprincess):

Actually i was typing part of it so u can do the remaining one. it is the same.:)

mathslover (mathslover):

Oh k . Thanks for letting me know.

OpenStudy (ajprincess):

Welcome:)